find sin (x),given that cos(x) = 1

Answer by stanbon(75887)

find sin (x),given that cos(x) = 1
---
cos(x) = 1 implies x = 1 and r = 1
--
Then y = sqrt(1^2 - 1^2] = 0
--
sin(x) = y/r = 0/1 = 0
=============================
Cheers,
Stan H.
=============